OpenStudy (anonymous):

Having some trouble figuring this out... If (x+7)/8 = (x+8)/9 Then what is x? Step by step explanations please.

OpenStudy (anonymous):

=>9(x+7)=8(x+8) expand:9x+63=8x+64 isolate x:9x-8x=64-63 =>x=1

OpenStudy (anonymous):

Thank you so much for your answer. Unfortunately when I was solving the problem I completely forgot to look at the denominators after I multiplied them... Not sure if that makes any sense. Anyhow, thank you so much!

OpenStudy (anonymous):

yw